The sum of three consecutive integers is 57. If n is the smallest integer, which equation could be used to determine n?
Answer by algebrahouse.com(1659)

n = first integer
n + 1 = second consecutive integer
n + 2 = third consecutive integer

n + n + 1 + n + 2 = 57 {sum of the three is 57}

For more help from me, visit: www.algebrahouse.com